OpenStudy (jusaquikie):

Find the derivative of the function. y = 5xe^−kx y'(x) =

OpenStudy (jusaquikie):

\[ y = 5xe^{-kx}\]

OpenStudy (jusaquikie):

i know to use the chain rule and the product rule but i don't know how to seperate this out properly

OpenStudy (anonymous):

y is made up of a product function. Thus, 5x and e^(-kx) Hence, we are to use the product rule with the above separations

OpenStudy (jusaquikie):

the answer has ln in it how do e and ln work together?

OpenStudy (anonymous):

Where from the In?

OpenStudy (anonymous):

If you differentiate y = e^ax, what do you get?

OpenStudy (jusaquikie):

it would be E^ax right?

OpenStudy (jusaquikie):

or would it be ae^ax-1

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = u \frac{ dv }{ dx } + v\frac{ du }{ dx }\]

OpenStudy (jusaquikie):

or (e^ax)a

OpenStudy (anonymous):

it will give you ae^ax

OpenStudy (jusaquikie):

constant letters confuse me more than anything

OpenStudy (anonymous):

Consider them as numbers

OpenStudy (jusaquikie):

but moving a to the front of the equason wouldn't i have a -1 on the exponent?

OpenStudy (anonymous):

no

OpenStudy (jusaquikie):

so 5xe^-kx is 5x*e^-kx so (5)(e^-kx)+(5x)(-ke^-kx) = 5e^-kx - 5xke^-kx

OpenStudy (anonymous):

So \[y=5xe^\left( -kx \right)\] \[\frac{ dy }{ dx } = 5x \frac{ d }{ dx }(e^{-kx}) + e ^{-kx} \frac{ d }{ dx }(5x)\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = 5x(-ke ^{-kx}) + e ^{-kx}(5)\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = -5kxe ^{-kx} + 5e ^{-kx}\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = -5e ^{-kx} (kx - 1)\]

OpenStudy (jusaquikie):

Thanks KKJ